# Learning Performance Maximizing Ensembles with Explainability Guarantees

### v2

@@ Introduction
In most high stakes settings, such as medical diagnosis (Gulum, Trombley, and Kantardzic 2021) and criminal justice (Rudin 2019), model predictions have two viability requirements. Firstly, they must exceed a given global performance threshold, thus ensuring an adequate understanding of the underlying process. Secondly, model predictions must be explainable.
Explainability, however defined (Linardatos, Papastefanopoulos, and Kotsiantis 2020), is a desirable characteristic in any prediction function. Intrinsically interpretable “glass box” models ((Agarwal et al. 2021), (Lemhadri, Ruan, and Tibshirani 2021), (Rymarczyk et al. 2020)), which are explainable by construction, are particularly advantageous as they require no additional post-hoc processing ((Ribeiro, Singh, and Guestrin 2016), (Lundberg and Lee 2017)) to achieve explainability, and thus also avoid complications arising from post-hoc explanation learning ((Rudin 2019), (Garreau and Luxburg 2020)). Due to these advantages, glass box models are uniquely suited to settings where faithful explanations of predictions are required.
@@
Figure 1: This figure shows a two-class classification task in which the areas of expertise (the diamond pattern for the glass box and the spiral pattern for the black box model) are complementary. In this example, the glass box achieves a $92.7\%$ accuracy, the black box reaches $95.0\%$ accuracy, and the allocated ensemble of the two exceeds both with a $95.8\%$ accuracy. Thus, the resulting EEG allocation improves performance over both component models while also providing explainability (for $20\%$ of observations in this case).
@@
Methods vary in the range of component models they can accommodate. Some are defined for only one set of glass box, black box, and allocator model types - for example LSP (Wang and Saligrama 2012) and OTSAM (Wang, Fujimaki, and Motohashi 2015), which use binary tree-type splitting to define regions, and linear models and sparse additive models respectively to predict within regions. Other methods are black box agnostic but still limited in glass box and allocator model type - for example HyRS (Wang 2019), HyPM (Wang and Lin 2021), CRL (Pan, Wang, and Hara 2020), and HybridCORELS (Ferry, Laberge, and Aïvodji 2023), which use rule-based models as both glass box and allocator. EEG is the only fully model-agnostic partial explainability method which can be implemented with any combination of glass box, black box, and allocator models.
@@
More extensive reviews of the partial explainability approach and explainability methods in general are available in (Linardatos, Papastefanopoulos, and Kotsiantis 2020), (Nauta et al. 2022), and (Sahakyan, Aung, and Rahwan 2021).
@@
Q-Complete Allocation Optimality: The final key point of novelty of the EEG approach is the optimality of allocation, as defined in Proposition 1 and Proposition 2, which is encoded in the allocation desirability ranking for any explainability level. Thus, the learned allocator, which estimates this ranking, is an explicit function of $q$ and provides the allocation solution to any explainability level after training only once. This capability is in contrast with previous works which provide, at most, several explainability level solutions with varying degrees of stability (Ferry, Laberge, and Aïvodji 2023).
@@
First, we define the underlying task as the estimation of the function $f(x)=y$ where $x\in\mathcal{X},y\in\mathcal{Y}$. We also define observations as $z=(x,y)\in\mathcal{X}\times\mathcal{Y}=\mathcal{Z}$, the training dataset $D^{n}=\{z_{i}:i\in\{1,...,n\},z_{i}\in\mathcal{Z}\}$, and loss function for the underlying task $l_{U}:\mathcal{Y}\times\mathcal{Y}\rightarrow\mathbb{R}$. Next, we define the ensemble component models - first, the intrinsically explainable glass box model as $g:\mathcal{X}\rightarrow\mathcal{Y}$ and the alternate, black box model as $b:\mathcal{X}\rightarrow\mathcal{Y}$, both of which are learned independently on the full training dataset $D^{n}$.
Next, we define the allocation task. We define the class of all allocator functions as $A=\{a:\mathcal{V}\rightarrow\{0,1\}\}$ and the class of all “proper” allocator functions as $\hat{A}=\{a:\hat{\mathcal{V}}\rightarrow\{0,1\}\}$, where $\mathcal{V}$ is a general space of inputs (typically $\mathcal{Z}$) and $\hat{\mathcal{V}}\subseteq\mathcal{V}$ containing only information available at allocation time. Next we define the class of $q$-explainable allocators as $A_{q}=\{a_{q}:a_{q}\in A,\frac{1}{n}\sum^{n}_{i=1}a_{q}(v_{i})=q\}$ and the corresponding class of “proper” $q$-explainable allocators as $\hat{A}_{q}=\{a_{q}:a_{q}\in\hat{A},\frac{1}{n}\sum^{n}_{i=1}a_{q}(v_{i})=q\}$, for $q\in\mathcal{Q}=\{\frac{i}{n}:i\in\{1,...,n\}\}\subseteq[0,1]$, with $q$ being the explainability level. Note, the set $A$ is used to define the optimal allocator, whereas the set $\hat{A}$ is searched to obtain an estimator of this optimum.
We next define indicators of performance sufficiency. These functions $s:\mathcal{Z}\rightarrow\{0,1\}$ should be thought of as context-dependent indicators of whether performance within a region of the feature space is sufficiently high to use the model in question reliably for explanation. For classification tasks, we define performance sufficiency as $s^{\prime}_{f}(z)=\mathbb{I}{\{f(x)=y\}}$, and for regression tasks as $s^{\prime}_{f}(z)=\mathbb{I}{\{l_{U}(f(x),y)<\epsilon\}}$, with $f:\mathcal{X}\rightarrow\mathcal{Y}$. In practice $\epsilon$ should be selected based on problem-specific context, however, lacking such context in the regression experiments conducted for this study, $\epsilon$ was selected to be the lower of the average validation losses of $g$ and $b$, as a reasonable threshold for prediction correctness. These sufficiency indicators generate the following partition of the data: $Z^{\prime}_{0}=\{z:z\in\mathcal{Z},s^{\prime}_{g}(z)+s^{\prime}_{b}(z)=0\},Z^{\prime}_{2}=\{z:z\in\mathcal{Z},s^{\prime}_{g}(z)+s^{\prime}_{b}(z)=2\},Z^{\prime}_{g}=\{z:z\in\mathcal{Z},s^{\prime}_{g}(z)=1,s^{\prime}_{b}(z)=0\}$, and $Z^{\prime}_{b}=\{z:z\in\mathcal{Z},s^{\prime}_{g}(z)=0,s^{\prime}_{b}(z)=1\}$, with $n_{0}=|Z^{\prime}_{0}|,n_{2}=|Z^{\prime}_{2}|,n_{g}=|Z^{\prime}_{g}|,n_{b}=|Z^{\prime}_{b}|$, and $n_{q}=nq$.
Next we motivate the use of the sufficiency perspective. Sufficiency functions are critical for defining coherent allocations when, as is often the case, the absolute performance measures used to learn ensemble component models do not match allocation preference (e.g. loss minimization vs accuracy maximization). Consider the following constrained allocation decision in which only one observation can be allocated to $g$:
Obs $l_{U}(g)$ $s_{g}$ $l_{U}(b)$ $s_{b}$
@@
In this case, loss minimization dictates an allocation of $z_{1}$ to $g$ and $z_{2}$ to $b$, which would allocate $z_{2}$ to an insufficient prediction. Sufficiency maximization would however yield a more satisfactory allocation of $z_{2}$ to $g$ and $z_{1}$ to $b$. This example demonstrates the utility of sufficiency allocation - distinguishing between a case where the user is willing to sacrifice “a bit of performance” (as quantified by sufficiency) for explanation ($z_{1}$), and a case were even a small performance drop results in an explanation that is not sufficiently trustworthy to use ($z_{2}$).
@@
In the allocation task, the objective is to construct an allocator $a_{q}$ that will determine which model, either the explainable $g$ or the black box $b$, is used for prediction on any given observation $z$, in a manner that is optimal relative to the following criteria. Firstly, for any given explainability level $q$, the allocator should distribute observations in a way that maximizes sufficient ensemble performance, defined as $\bar{t}^{\prime}(a_{q})=\frac{1}{n}\sum_{i=1}^{n}s_{g}^{\prime}(z_{i})a_{q}(v_{i})+s_{b}^{\prime}(z_{i})(1-a_{q}(v_{i}))$. Secondly, and again for any $q$, the allocator should maximize sufficient explainable prediction $\bar{t}^{\prime}_{g}(a_{q})=\frac{1}{n}\sum_{i=1}^{n}s_{g}^{\prime}(z_{i})a_{q}(v_{i})$, i.e. the performance of the model $g$ on the subset of observations it has been allocated, subject to maintaining maximal $\bar{t}^{\prime}(a_{q})$. Finally, the allocator should also be consistent in its allocations across the values of $q$, meaning that if an observation is allocated to $g$ for a given $q$, it should remain allocated to the glass box for all higher explainability levels as well. Next, we define our allocator, and show that it meets the three criteria introduced above.
Our proposed allocation function is defined as $a^{\prime}_{q}(z)=\mathbb{I}\{r^{\prime}(z)>1-q\}$, where rescaled ranking $r^{\prime}(z)=\frac{\mathrm{rank}_{D^{n}}(\tilde{r}^{\prime}(z))}{n}$, and ranking $\tilde{r}^{\prime}(z)=2s^{\prime}_{g}(z)-s^{\prime}_{b}(z)-\sigma(l_{U}(b(x),y)-l_{U}(g(x),y))$, with $\sigma(x)=\frac{1}{1+e^{-x}}$.
@@
The intuition behind the allocator is as follows. First, all observations are sorted in sufficient performance maximizing order, i.e. allocation of observations in $Z_{g}^{\prime}$ to $g$ is prioritized over allocation of observations in $Z_{2}^{\prime}$ and $Z_{0}^{\prime}$, which in turn are prioritized over $Z_{b}^{\prime}$. Next, observations are sorted in explainable sufficient performance maximizing order, i.e. $Z_{2}^{\prime}$ is prioritized ahead of $Z_{0}^{\prime}$ for allocation to $g$. Then, within each sufficiency category, observations are ordered in absolute performance maximizing order, i.e. observations with large relative performance of $g$ over $b$ are prioritized for allocation to $g$. Next, this ranking is normalized, yielding the glass box allocation desirability percentile $r^{\prime}$. An important feature of this percentile is that it is constant with respect to $q$, thus the optimal observations to allocate to $g$, for any level of $q$, are simply the $n_{q}$ most highly ranked, resulting in the allocator $a^{\prime}_{q}$.
@@
Note that in the described methodology, sufficiency based allocation can be viewed as a generalization of allocation via absolute performance, and can thus be reduced to the latter by selecting either $s_{f}(z)=0$ or $s_{f}(z)=1$, $\forall z,f$.
@@
Next, we state the optimality properties of the proposed allocator $a^{\prime}_{q}$. The proofs are available in the long form paper on arxiv.org in the Theoretical Results section of the Appendix.
@@
(Maximal Sufficient Performance) $\forall q\in\mathcal{Q},a^{\prime}_{q}\in A^{*}_{q}$ where $A^{*}_{q}=\{a^{*}_{q}\colon a^{*}_{q}=\arg\max_{a_{q}\in A_{q}}\bar{t}^{\prime}(a_{q})\}$ and $\bar{t}^{\prime}(a_{q})=\frac{1}{n}\sum_{i=1}^{n}t^{\prime}(a_{q},z_{i})=\frac{1}{n}\sum_{i=1}^{n}s^{\prime}_{g}(z_{i})a_{q}(z_{i})+s^{\prime}_{b}(z_{i})(1-a_{q}(z_{i}))$
@@
(Maximal Sufficient Explainable Performance) $\forall q\in\mathcal{Q},a^{\prime}_{q}\in A^{*}_{q|g}$ where $A^{*}_{q|g}=\{a^{*}_{q|g}\colon a^{*}_{q|g}=\arg\max_{a^{*}_{q}\in A^{*}_{q}}\bar{t}^{\prime}_{g}(a^{*}_{q})\}$ and $\bar{t}^{\prime}_{g}(a^{*}_{q})=\frac{1}{n}\sum_{i=1}^{n}t^{\prime}_{g}(a^{*}_{q},z_{i})=\frac{1}{n}\sum_{i=1}^{n}s^{\prime}_{g}(z_{i})a^{*}_{q}(z_{i})$
@@
(Monotone Allocation) $\forall q_{i}<q_{j}\in\mathcal{Q}$, $\{z:z\in\mathcal{Z},a^{\prime}_{q_{i}}(z)=1\}\subseteq\{z:z\in\mathcal{Z},a^{\prime}_{q_{j}}(z)=1\}$
@@
Tabular data is used to evaluate the proposed methodology as it the setting for which the required intrinsically explainable glass box models are most readily available. Following the tabular data benchmarking framework proposed by (Grinsztajn, Oyallon, and Varoquaux 2022), we conduct experiments on a set of 31 datasets (13 classification, 18 regression). These datasets represent the full set of provided datasets with quantitative features less the four largest scale datasets (omitted due to computational limitations). These datasets are summarized in Table 4.
Each dataset is split (70%, 9%, 21%) into training, validation, and test sets respectively, following (Grinsztajn, Oyallon, and Varoquaux 2022). All features and regression response variables are rescaled to the range [-1,1].
@@
For allocator training, the features $x$ are augmented with four additional constructed features, the predictions $g(x)$ and $b(x)$, and two distance measures $d(g(x),b(x))$ between them, the cross-entropy and MSE. In our experiments, inclusion of these features improved allocator learning - likely by removing the need for the allocator to attempt to learn these quantities on its own. Allocation performance is further improved by ensembling the feature-dependent learned allocator $a^{\prime}_{q}$ with a strong feature-independent allocator $a^{\prime\prime}_{q}$, where $a^{\prime\prime}_{q}(d(g(z),b(z)))=\mathbb{I}\{\frac{\mathrm{rank}_{D^{n}}d(g(z),b(z))}{n}>1-q\}$. $a^{\prime\prime}_{q}$ can be viewed as an “assume the black box is correct” allocation rule which is more likely to assign an observation to $g$ if the distance between the predictions of $g$ and $b$ is high. Which of the two allocators is used for a given $q$ is determined by their respective performances on the validation set.
@@
Hyperparameter tuning for all models is done using 4-fold cross-validation, with the exception of the neural network tuning which is done using the validation set. A grid search is done to select the best hyperparameters for each model with search values available in Tables 5, 6, and 7.
@@
Finally, we define the accuracy with which the four sufficiency categories ($Z^{\prime}_{g}$, $Z^{\prime}_{b}$, $Z^{\prime}_{2}$, and $Z^{\prime}_{0}$) can be estimated as the sufficiency accuracy ($s$ Acc). The higher this accuracy, the better able the allocator is to inform the user of which category a given observation is likely to be a member of.

### v3

@@ Introduction
In most high stakes settings, such as medical diagnosis (Gulum et al. 2021) and criminal justice (Rudin 2019), model predictions have two viability requirements. Firstly, they must exceed a given global performance threshold, thus ensuring an adequate understanding of the underlying process. Secondly, model predictions must be explainable.
Explainability, however defined (Linardatos et al. 2020), is a desirable characteristic in any prediction function. Intrinsically interpretable “glass box” models ((Agarwal et al. 2021), (Lemhadri et al. 2021), (Rymarczyk et al. 2020)), which are explainable by construction, are particularly advantageous as they require no additional post-hoc processing ((Ribeiro et al. 2016), (Lundberg and Lee 2017)) to achieve explainability, and thus also avoid complications arising from post-hoc explanation learning ((Rudin 2019), (Garreau and Luxburg 2020)). Due to these advantages, glass box models are uniquely suited to settings where faithful explanations of predictions are required.
@@
Figure 1: This figure shows a two-class classification task in which the areas of expertise (the diamond pattern for the glass box and the spiral pattern for the black box model) are complementary. The glass box achieves a $92.7\%$ accuracy, the black box reaches $95.0\%$ accuracy, and the allocated ensemble of the two exceeds both with a $95.8\%$ accuracy. Thus, the resulting EEG allocation improves performance over both component models while also providing explainability (for $20\%$ of observations in this case).
Figure 2: This figure provides an intuition for how the proposed allocator $a^{\prime}_{q}(v)$ ranks observations. This ranking can be seen as ordering the sufficiency sets ($Z_{b},Z_{0},Z_{2},Z_{g}$) step-by-step, with each step achieving a particular optimality condition (formalized in Propositions 1, 2, and 3). In the figure, individual observations are circles above the number line with size representing the magnitude of the relative advantage of black box over glass box, the left-hand color representing sufficiency of the glass box, and the right-hand color representing the sufficiency of the black box (green for sufficient, red for insufficient).
@@
Methods vary in the range of component models they can accommodate. Some are defined for only one set of glass box, black box, and allocator model types - for example LSP (Wang and Saligrama 2012) and OTSAM (Wang et al. 2015), which use binary tree-type splitting to define regions, and linear models and sparse additive models respectively to predict within regions. Other methods are black box agnostic but still limited in glass box and allocator model type - for example HyRS (Wang 2019), HyPM (Wang and Lin 2021), CRL (Pan et al. 2020), and HybridCORELS (Ferry et al. 2023), which use rule-based models as both glass box and allocator. EEG is the only fully model-agnostic partial explainability method which can be implemented with any combination of glass box, black box, and allocator models.
@@
More extensive reviews of the partial explainability approach and explainability methods in general are available in (Linardatos et al. 2020), (Nauta et al. 2022), and (Sahakyan et al. 2021).
@@
Q-Complete Allocation Optimality: The final key point of novelty of the EEG approach is the optimality of allocation, as defined in Propositions 1, 2, 3, and 4, which is encoded in the allocation desirability ranking for any explainability level. Thus, the learned allocator, which estimates this ranking, is an explicit function of $q$ and provides the allocation solution to any explainability level after training only once. This capability is in contrast with previous works which provide, at most, several explainability level solutions with varying degrees of stability (Ferry et al. 2023).
@@
First, we define the underlying task as the estimation of the function $f(x)=y,f\in\mathcal{F}=\{f:\mathcal{X}\rightarrow\mathcal{Y}\}$ where $x\in\mathcal{X},y\in\mathcal{Y}$. We also define observations as $z=(x,y)\in\mathcal{X}\times\mathcal{Y}=\mathcal{Z}$, the training dataset $D^{n}=\{z_{i}:i\in\{1,...,n\},z_{i}\in\mathcal{Z}\}$, and loss function for the underlying task $l:\mathcal{Y}\times\mathcal{Y}\rightarrow\mathbb{R}$. Next, we define the ensemble component models - first, the intrinsically explainable glass box model as $g:\mathcal{X}\rightarrow\mathcal{Y}$ and the alternate, black box model as $b:\mathcal{X}\rightarrow\mathcal{Y}$, both of which are learned independently on the full training dataset $D^{n}$.
[…]
We next define indicators of performance sufficiency. These functions $s:\mathcal{F}\times\mathcal{Z}\rightarrow\{0,1\}$ should be thought of as context-dependent indicators of whether performance within a region of the feature space is sufficiently high to use the model in question reliably for explanation. Although the EEG approach holds for any such function $s$, sufficiency functions used in the Experiments section are defined as follows. For classification tasks, we define performance sufficiency as $s_{f}(z)=\mathbb{I}{\{f(x)=y\}}$, and for regression tasks as $s_{f}(z)=\mathbb{I}{\{l(f(x),y)<\epsilon\}}$. In practice $\epsilon$ should be selected based on problem-specific context, however, lacking such context in the regression experiments conducted for this study, $\epsilon$ was selected to be the lower of the average validation losses of $g$ and $b$, as a reasonable threshold for prediction correctness. These sufficiency indicators generate the following partition of the data: $Z_{0}=\{z:z\in\mathcal{Z},s_{g}(z)+s_{b}(z)=0\},Z_{2}=\{z:z\in\mathcal{Z},s_{g}(z)+s_{b}(z)=2\},Z_{g}=\{z:z\in\mathcal{Z},s_{g}(z)=1,s_{b}(z)=0\}$, and $Z_{b}=\{z:z\in\mathcal{Z},s_{g}(z)=0,s_{b}(z)=1\}$, with $n_{0}=|Z_{0}|,n_{2}=|Z_{2}|,n_{g}=|Z_{g}|,n_{b}=|Z_{b}|$, and $n_{q}=nq$.
Next we motivate the use of the sufficiency perspective. Sufficiency functions are critical for defining coherent allocations when, as is often the case, the absolute performance measures used to learn ensemble component models do not match allocation preference (e.g. loss minimization vs accuracy maximization). Consider the constrained allocation decision in Table 1, in which only one observation can be allocated to $g$.
Obs $l(g)$ $s_{g}$ $l(b)$ $s_{b}$
@@
Table 1: This table describes the loss values ($l$) and sufficiencies ($s$) of two observations ($z_{1}$, $z_{2}$), for both a glass box ($g$) and black box model ($b$). In the constrained allocation case, in which only one observation can be allocated to $g$, the optimal allocation changes depending on whether loss or sufficiency is used to determine allocation preference.
In this case, loss minimization dictates an allocation of $z_{1}$ to $g$ and $z_{2}$ to $b$, which would allocate $z_{2}$ to an insufficient prediction. Sufficiency maximization would however yield a more satisfactory allocation of $z_{2}$ to $g$ and $z_{1}$ to $b$. This example demonstrates the utility of sufficiency allocation - distinguishing between a case where the user is willing to sacrifice “a bit of performance” (as quantified by sufficiency) for explanation ($z_{1}$), and a case where even a small performance drop results in an explanation that is not sufficiently trustworthy to use ($z_{2}$).
@@
In the allocation task, the objective is to construct an allocator $a_{q}$ that will determine which model, either the explainable $g$ or the black box $b$, is used for prediction on any given observation $z$, in a manner that is optimal relative to the following criteria. Firstly, for any given explainability level $q$, the allocator should distribute observations in a way that maximizes sufficient ensemble performance, defined as
$S(a_{q})=\frac{1}{n}\sum^{n}_{i=1}s(a_{q},v_{i})$ (1)
$s(a_{q},v)=s_{g}(z)a_{q}(v)+s_{b}(z)(1-a_{q}(v))$ (2)
Secondly, and again for any $q$, the allocator should maximize sufficient explainable prediction, defined as
$S_{g}(a_{q})=\frac{1}{n}\sum^{n}_{i=1}s_{g}(z_{i})a_{q}(v_{i})$ (3)
i.e. the performance of the model $g$ on the subset of observations it has been allocated. Thirdly, for any $q$, the allocator should maximize the ensemble absolute performance (as measured by $L(a_{q})$ below), conditional on achieving the first two sufficient performance objectives.
$\begin{split}L(a_{q}&)=\frac{1}{n}\sum l_{i}(a_{q})\\ =\frac{1}{n}\sum l(g(x_{i}),y_{i})a_{q}(&v_{i})+l(b(x_{i}),y_{i})(1-a_{q}(v_{i}))\end{split}$ (4)
Finally, the allocator should also be consistent in its allocations across the values of $q$, meaning that if an observation is allocated to $g$ for a given $q$, it should remain allocated to the glass box for all higher explainability levels as well. Next, we define our allocator, and show that it meets the criteria introduced above.
Our proposed allocation function $a^{\prime}_{q}(v)$ is defined as follows.
$a^{\prime}_{q}(v)=\mathbb{I}\{r(z)>1-q\}$ (5)
With rescaled ranking $\tilde{r}$ defined as
$\tilde{r}(z)=\frac{\mathrm{rank}_{D^{n}}(r(z))}{n}$ (6)
and ranking $r$ defined as
[…]
where $\sigma(x)=\frac{1}{1+e^{-x}}$.
@@
The intuition behind the allocator is depicted in Fig. 2 and is as follows. First, all observations are sorted in sufficient performance maximizing order, i.e. allocation of observations in $Z_{g}$ to $g$ is prioritized over allocation of observations in $Z_{2}$ and $Z_{0}$, which in turn are prioritized over $Z_{b}$. Next, observations are sorted in explainable sufficient performance maximizing order, i.e. $Z_{2}$ is prioritized ahead of $Z_{0}$ for allocation to $g$. Then, within each sufficiency category, observations are ordered in absolute performance maximizing order, i.e. observations with large relative performance of $g$ over $b$ are prioritized for allocation to $g$. Next, this ranking is normalized, yielding the glass box allocation desirability percentile $r$. An important feature of this percentile is that it is constant with respect to $q$, thus the optimal observations to allocate to $g$, for any level of $q$, are simply the $n_{q}$ most highly ranked, resulting in the allocator $a^{\prime}_{q}$.
@@
Note that in the described methodology, sufficiency based allocation can be viewed as a generalization of allocation via absolute performance, and can thus be reduced to the latter by selecting, for example, either $s_{f}(z)=0$ or $s_{f}(z)=1$, $\forall z,f$.
@@
Next, we state the optimality properties of the proposed allocator $a^{\prime}_{q}$. The proofs are available in the Theoretical Results section of the Appendix.
@@
(Maximal Sufficient Performance)
$\forall q\in\mathcal{Q},a^{\prime}_{q}\in A^{*}_{q}$ where $A^{*}_{q}=\{a^{*}_{q}\colon a^{*}_{q}=\arg\max_{a_{q}\in A_{q}}S(a_{q})\}$
In words, Proposition 1 states that across all explainability levels $q$, the proposed allocator $a^{\prime}_{q}(v)$ (Eq. 5) achieves the highest possible sufficient performance (as defined in Eq. 1).
@@
(Maximal Sufficient Explainable Performance)
$\forall q\in\mathcal{Q},a^{\prime}_{q}\in A^{*}_{q|g}$ where $A^{*}_{q|g}=\{a^{*}_{q|g}\colon a^{*}_{q|g}=\arg\max_{a_{q}\in A_{q}}S_{g}(a_{q})\}$
Proposition 2 states that, again across all explainability levels, the proposed $a^{\prime}_{q}(v)$ achieves the highest possible sufficient performance on those observations which have been allocated to the glass box $g$ (i.e. maximizing Eq. 3).
@@
(Conditional Maximum Absolute Performance)
$\forall q\in\mathcal{Q},a^{\prime}_{q}\in A^{*}_{q|s}$ where $A^{*}_{q|s}=\{a^{*}_{q|s}\colon a^{*}_{q|s}=\arg\min_{a_{q}\in A^{*}_{q}\cap A^{*}_{q|g}}L(a_{q})\}$
Proposition 3 states that across all explainability levels, the proposed allocator $a^{\prime}_{q}(v)$ achieves the highest absolute performance (in the sense of Eq. 4) available among all allocators that meet the conditions of Prop 1 and Prop 2.
Proposition 4.
(Monotone Allocation)
$\forall q_{i}<q_{j}\in\mathcal{Q}$, $\{v:v\in\mathcal{V},a^{\prime}_{q_{i}}(v)=1\}\subseteq\{v:v\in\mathcal{V},a^{\prime}_{q_{j}}(v)=1\}$
Finally, Proposition 4 states that the allocator $a^{\prime}_{q}(v)$ is self-consistent in the sense that if an observation is considered best allocated to the glass box at a given explainability level $q_{i}$, then is also guaranteed to remain allocated to the glass box at any higher higher explainability level $q_{j}$. This prevents the unintuitive phenomenon of an observation ”hopping” in and out of the glass-box allocated bucket at different explainability levels.
@@
Tabular data is used to evaluate the proposed methodology as it the setting for which the required intrinsically explainable glass box models are most readily available. Following the tabular data benchmarking framework proposed by (Grinsztajn et al. 2022), we conduct experiments on a set of 31 datasets (13 classification, 18 regression). These datasets represent the full set of provided datasets with quantitative features less the four largest scale datasets (omitted due to computational limitations). These datasets are summarized in Section B of the Appendix.
Each dataset is split (70%, 9%, 21%) into training, validation, and test sets respectively, following (Grinsztajn et al. 2022). All features and regression response variables are rescaled to the range [-1,1].
@@
For allocator training, the features $x$ are augmented with four additional constructed features, the predictions $g(x)$ and $b(x)$, and two distance measures $d(g(x),b(x))$ between them, the cross-entropy and MSE. In our experiments, inclusion of these features improved allocator learning - likely by removing the need for the allocator to attempt to learn these quantities on its own. Allocation performance is further improved by ensembling the feature-dependent learned allocator $a^{\prime}_{q}$ with a strong feature-independent allocator $a^{\prime\prime}_{q}$, where $a^{\prime\prime}_{q}(d(g(z),b(z)))=\mathbb{I}\{\frac{\mathrm{rank}_{D^{n}}d(g(z),b(z))}{n}<q\}$. $a^{\prime\prime}_{q}$ can be viewed as an “assume the black box is correct” allocation rule which is more likely to assign an observation to $g$ if the distance between the predictions of $g$ and $b$ is low. Which of the two allocators is used for a given $q$ is determined by their respective performances on the validation set.
@@
Hyperparameter tuning for all models is done using 4-fold cross-validation, with the exception of the neural network tuning which is done using the validation set. A grid search is done to select the best hyperparameters for each model with search values available in Section C of the Appendix.
@@
Finally, we define the accuracy with which the four sufficiency categories ($Z_{g}$, $Z_{b}$, $Z_{2}$, and $Z_{0}$) can be estimated as the sufficiency accuracy ($s$ Acc). The higher this accuracy, the better able the allocator is to inform the user of which category a given observation is likely to be a member of.